OpenStudy (anonymous):

the line containing (2,-4) and (5,1)

OpenStudy (anonymous):

1+4 over5-2=5/3 is the slope of the line and then you can start at -4 so the line equation y=5/3x-4

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

I hope that is what you were looking for

OpenStudy (anonymous):

it was thank you very much :))